A pharmacist mixed some 5%-saline solution with some 11%-saline solution to obtain 100 mL of a 9% solution. How much of the 5% solution did the pharmacist use in the mixture?

Answer by mananth(16946)

percent ---------------- quantity
Saline solution I 5.00% ---------------- x ml
Saline solution II 11.00% ------ 100 - x ml
Mixture 9.00% ---------------- 100
12
5.00% x + 11.00% ( 100 - x ) = 9.00% * 100
5 x + 11 ( 100 - x ) = 900
5 x + 1100 - 11 x = 900
5 x - 11 x = 900 - -1100
-6 x = -200
/ -6
x = 33.33 ml 5.00% Saline solution I
66.67 ml 11.00% Saline solution II

m.ananth@hotmail.ca